A circle passes through the points (-3,4),(1,0)and its centre lies on the x-axis .Find the equation of the circle

Grade:Select Grade

1 Answers

Vikas TU
6 years ago
since center lies on the x axis.The center can be represented by (h,0)
The distance between (h,0) and (1,0) is radius since both lie on the x axis.
And circle passing through (-3,4) can be represented by
(-3 –h)2 + (4 – 0)2 = r2
=> 9 + h^2 + 6h + 16 = r^2-----------------1
 circle passing through (1,0) can be represented by
(1-h)2 = r------------------------2
Equating 1 and 2
9 + h^2 + 6h + 16 = 1 + h^2 – 2h
8h = -24
h = -3.
Radius is ((1-h)2 )^ ½
Radius is 4